What are the vertices of the ellipse whose equation is (x2)/9 + (y2)/36 = 1?
Answer by lwsshak3(11628)

What are the vertices of the ellipse whose equation is (x2)/9 + (y2)/36 = 1
**
Standard form of equation for ellipse with vertical major axis (a^2 under y^2):
(x-h)^2/b^2+(y-k)^2/a^2=1,a>b, (h,k) being the (x,y) coordinates of the center.
For given equation: x^2/9+y^2/36=1
center: (0,0)
a^2=36
a=√36=6
length of vertical major axis=2a=12
Vertices are end points of major axis=(0,0±a)=(0,0±6)=(0,-6) and (0,6)
Vertices are at (0,-6) and (0,6)